OpenStudy (anonymous):

find the coordinate of the points which divides the line segment from (-2,1) and (2,-3) in the ratio 3 to 4

OpenStudy (anonymous):

can you ???

ganeshie8 (ganeshie8):

use the section formula : point which divides the line segment from \((x_1, y_1)\) and \((x_2, y_2)\) in ratio \(m : n\) is \[\left( \dfrac{mx_2 + nx_1}{m+n}, \dfrac{my_2 + ny_1}{m+n}\right)\]

ganeshie8 (ganeshie8):

simply plugin the values

OpenStudy (anonymous):

and sir ??

ganeshie8 (ganeshie8):

\[\left( \dfrac{3(2) + 4(-2)}{3+4},~ \dfrac{3(-3) + 4(1)}{3+4}\right)\]

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

-2over1 and -5over7 ???

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

\((-\frac{2}{7}, -\frac{5}{7})\) is the required point

OpenStudy (anonymous):

thank you now i know

ganeshie8 (ganeshie8):

u wlc :)